ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 00A, 26A Bilotskiyi Nikolay N., National Pedagodic University, Ukraine Khilchenko, Luydmila Y., Yavneh Academy, USA An algorithmic approach to the notion of an elementary function An elementary function is one of the basics notion of precalculus and calculus courses. But many calculus books do not include the definition or are abounded with many errors, connected with the idea of an elementary function. The analysis of these errors could be found in [1]. But the suggested correct definition of an elementary function is too complicated for teaching to beginners. We offer a new rigorous and understandable to students approach to the notion of an elementary function based on the teaching to beginners. We offer a new rigorous and understandable to students approach to the notion of an elementary function based on the concept of linear algorithm. Part of these results is a joint work with I. Subbotin. Reference: B. P. Pygachov, On some error in calculus manuals, Uspehi. Math. Nayk., 19 (1964), 5, 234.